A drawer contains 50 bolts & 150 nuts ,half of the bolts are rusted,if one item is choosen at random .what is the probability that it is rusted or a bolt?

1
Expert's answer
2021-12-31T04:35:54-0500

Total bolts = 50

Total nuts = 150

Total number of items in the bag = 50 + 150 = 200

Total number of rusted bolts = 50 / 2 = 25

If one item is chosen at random then the probability that it is bolt will be:


"P_1=\frac{50}{200}=\frac{1}{4}"


If one item is chosen at random then the probability that it is rusted will be:


"P_2=\frac{25}{200}=\frac{1}{8}"


The probability that randomly selected item is rusted or a bolt:


"P=P_1+P_2"


"P=\frac{1}{4}+\frac{1}{8}"


"P=\frac{3}{8}\approx0.375"